OpenStudy (anonymous):

An online movie rental store rented out 6,816 movies one month. This was 40% less than the number of movies it rented out the previous month. The following equation represents this situation, where x represents the number of movies the store rented out the previous month. x − 0.4x = 6,816 How many movies did the store rent out in both months combined?

OpenStudy (anonymous):

solve for x x − 0.4x = 6,816 x=11360 add 11360 and 6,816 to get the total

OpenStudy (anonymous):

From the equation 0.6x = 6816 x = 6816/0.6 = 11360 The store rented out x movies in a month and 40% less (OR 60% of x) the next month. That means that in both months combined it rented out 160% of x movies. So it rented out (160/100)*x = (160/100)*11360 = 18176 18,176 movies in both months combined